A glass bottle full of mercury has a mass of 500g being heated through 35 degree celsius. 12.43 g of mercury was expelled. Calculate the mass of mercury left in the bottle (cubic expansivity of mercury 1.8 X 10^-4, linear expansivity of glass is 8.0 X10^-6)
1
Expert's answer
2019-06-30T18:52:47-0400
"\Delta m=m-m'"

The mass of mercury left in the bottle:

"m'=m-\Delta m"

"m'=500-12.43=487.57\ g"